# Arithmetic of cuts in ordered abelian groups and of ideals over valuation rings

### v4

@@ 3. $\mathcal{O}$-ideals
We consider (possibly fractional) $\mathcal{O}_{v}$-ideals $I$, $I_{1}\,$, $I_{2}\,$, $J$, etc. We are interested in the inequality
@@ 3.1. Overrings of $\mathcal{O}_{v}$ and their ideals
[…]
@@
(43) $H(\mathcal{O})\>=\>{\mathcal{G}}(v\mathcal{O})\>=\>{\mathcal{G}}(v\mathcal{M})\;\mbox{\ \ and\ \ }\;\mathcal{O}\>=\>\mathcal{O}({\mathcal{G}}(v\mathcal{O}))\>=\>\mathcal{O}({\mathcal{G}}(v\mathcal{M}))\>.$
@@ 3.2. The invariance valuation ring of an $\mathcal{O}_{v}$-ideal
(45) $\mathcal{O}(\mathcal{O})\>=\>\mathcal{O}(\mathcal{M})\>=\>\mathcal{O}\;\;\;\mbox{\ \ and\ \ }\;\;\;\mathcal{M}(\mathcal{O})\>=\>\mathcal{M}(\mathcal{M})\>=\>\mathcal{M}\>.$
@@ 3.3. Properties of $\mathcal{M}(I)$
For $\mathcal{O}_{v}$-ideals $I$, the ideals $\mathcal{M}(I)$ appear in [7, Chapter II, Section 4] under the notation $I^{\sharp}$. There, properties of $\mathcal{M}(I)=I^{\sharp}$ are listed. In what follows, we will present several of them and show how they follow from the work we have done so far, in particular from our results on final segments in ordered abelien group in Section 2.
@@ 3.4. The equation $I_{1}J=I_{2}$
$\widehat{J}_{\mathcal{O}}\>:=\>\left\{\begin{array}[]{ll}a\mathcal{O}&\mbox{if there exists $a\in K$ such that $wa$ is the infimum of $wJ$, and}\\ J&\mbox{otherwise.}\end{array}\right.$
@@
()(*)( ∗ )  $\mathcal{O}(I_{1})\subseteq\mathcal{O}(I_{2})$, and $I_{2}$ is a nonprincipal $\mathcal{O}(I_{2})$-ideal if $I_{1}$ is a nonprincipal $\mathcal{O}(I_{2})$-ideal
@@
If condition ()(*)( ∗ ) does not hold, or equivalently, one of the cases
@@
Since $\mathcal{O}(I)=\mathcal{O}({\mathcal{G}}(vI))$ for every $\mathcal{O}_{v}$-ideal $I$ by equation (46), we know that $\mathcal{O}(I_{1})\subseteq\mathcal{O}(I_{2})$ if and only if ${\mathcal{G}}(vI_{1})\subseteq{\mathcal{G}}(vI_{2})$. Part 2) of Lemma 3.2, with the valuation $w$ associated with $\mathcal{O}(I_{2})$ in place of $v$, shows for $i=1,2$ that $I_{i}$ is a principal $\mathcal{O}(I_{2})$-ideal if and only if $wI_{i}=vI_{i}/{\mathcal{G}}(I_{2})$ is principal. Further, if $vI_{1}/{\mathcal{G}}(I_{2})$ is nonprincipal, then also $I_{1}$ is nonprincipal and therefore, $vI_{2}(\mathcal{O}_{v}:I_{1})=vI_{2}-(vI_{1})^{c}$ by (36).
@@ 3.5. Annihilators of quotients $I_{1}/I_{2}$
$IJ:I=\>\left\{\begin{array}[]{ll}J\mathcal{O}(IJ)&\mbox{if $I\mathcal{O}(IJ)$ is a principal $\mathcal{O}(IJ)$-ideal, and}\\ (J\mathcal{O}(IJ))^{\diamondsuit}&\mbox{if $I\mathcal{O}(IJ)$ is a nonprincipal $\mathcal{O}(IJ)$-ideal.}\end{array}\right.$
@@
2) Assume that (6) does not have a solution, and let $I^{\prime}_{2}$ be as defined in Theorem 3.14. Then the annihilator of $I_{1}/I_{2}$ is $a^{-1}I^{\prime}_{2}$ if $I^{\prime}_{1}\mathcal{O}(I_{2})=a\mathcal{O}(I_{2})$, and is $I^{\prime}_{2}(\mathcal{O}_{v}:I_{1})^{\diamondsuit}$ if $I_{1}\mathcal{O}(I_{2})$ is a nonprincipal $\mathcal{O}(I_{2})$-ideal.
@@ 3.6. Application to the computation of annihilators in special cases
Since $vJ=vb+(n-1)vI$, part 3) of Lemma 2.13 together with Corollary 2.14 shows that ${\mathcal{G}}(vI)={\mathcal{G}}(vJ)$, which implies that $\mathcal{O}(I)=\mathcal{O}(J)$.
@@
[…]
@@
$\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}\>=\>\left\{\begin{array}[]{ll}(a\mathcal{M})^{n-1}&\mbox{if $\mathcal{M}$ is a principal $\mathcal{O}$-ideal, }\\ (a\mathcal{O})^{n-1}=a^{n-1}\mathcal{O}&\mbox{if $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal.}\end{array}\right.$
@@
2): Assume first that $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}=\mathcal{M}_{v}\,$. Then $a\mathcal{M}/(a\mathcal{M})^{n}$ is nontrivial, whence $a\mathcal{M}\neq(a\mathcal{M})^{n}$, which by part 3) of Lemma 3.20 means that $a\in\mathcal{M}$ or $\mathcal{M}$ is a principal $\mathcal{O}$-ideal. If the latter holds, then by part 1), $a\mathcal{M}:(a\mathcal{M})^{n}=(a\mathcal{M})^{n-1}$. Since $a\mathcal{M}\subseteq\mathcal{M}_{v}\,$, this is equal to $\mathcal{M}_{v}$ if and only if $n-1=1$, $\mathcal{M}=\mathcal{M}_{v}\,$, and $va=0$, where we have used parts 1) and 3) of Lemma 3.20. On the other hand, if the latter does not hold, then $a\in\mathcal{M}$ and by part 1), $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}=(a\mathcal{O})^{n-1}\subseteq a\mathcal{O}$. The latter is contained in $\mathcal{M}$ but is not equal to $\mathcal{M}$ since $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal. Thus $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}=a\mathcal{M}:(a\mathcal{M})^{n}\subsetneq\mathcal{M}\subseteq\mathcal{M}_{v}$ and consequently, $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}\neq\mathcal{M}_{v}\,$.

### v5

@@ 3. $\mathcal{O}$-ideals
Recall that by an $\mathcal{O}_{v}$-ideal $I$ we mean a nonzero $\mathcal{O}_{v}$-module $I\subsetneq K$. We consider (possibly fractional) $\mathcal{O}_{v}$-ideals $I$, $I_{1}\,$, $I_{2}\,$, $J$, etc. We are interested in the inequality
@@ 3.1. Overrings of $\mathcal{O}_{v}$ and their ideals
This is a convex subgroup of the value group $vK$ of $(K,v)$. The valuation $w$ associated with $\mathcal{O}$ is (up to equivalence) given by
@@
[…]
@@ 3.2. The invariance valuation ring of an $\mathcal{O}_{v}$-ideal
[…]
@@ 3.3. Properties of $\mathcal{M}(I)$
For $\mathcal{O}_{v}$-ideals $I$, the ideals $\mathcal{M}(I)$ appear in [7, Chapter II, Section 4] under the notation $I^{\sharp}$. There, properties of $\mathcal{M}(I)=I^{\sharp}$ are listed. In what follows, we will present several of them and show how they follow from the work we have done so far, in particular from our results on final segments in ordered abelian groups in Section 2.
@@ 3.4. The equation $I_{1}J=I_{2}$
[…]
@@
()(*) $\mathcal{O}(I_{1})\subseteq\mathcal{O}(I_{2})$, and $I_{2}$ is a nonprincipal $\mathcal{O}(I_{2})$-ideal if $I_{1}$ is a nonprincipal $\mathcal{O}(I_{2})$-ideal
@@
If condition ()(*) does not hold, or equivalently, one of the cases
@@
Since $\mathcal{O}(I)=\mathcal{O}({\mathcal{G}}(vI))$ for every $\mathcal{O}_{v}$-ideal II by equation (46), we know that $\mathcal{O}(I_{1})\subseteq\mathcal{O}(I_{2})$ if and only if ${\mathcal{G}}(vI_{1})\subseteq{\mathcal{G}}(vI_{2})$. Part 2) of Lemma 3.2, with the valuation $w$ associated with $\mathcal{O}(I_{2})$ in place of $v$, shows for $i=1,2$ that $I_{i}$ is a principal $\mathcal{O}(I_{2})$-ideal if and only if $wI_{i}=vI_{i}/{\mathcal{G}}(I_{2})$ is principal. Further, if $vI_{1}/{\mathcal{G}}(I_{2})$ is nonprincipal, then also $I_{1}$ is nonprincipal and therefore, $vI_{2}(\mathcal{O}_{v}:I_{1})=vI_{2}-(vI_{1})^{c}$ by (36).
@@ 3.5. Annihilators of quotients $I_{1}/I_{2}$
[…]
@@
2) Assume that (6) does not have a solution, and let $I^{\prime}_{2}$ be as defined in Theorem 3.14. Then the annihilator of $I_{1}/I_{2}$ is $a^{-1}I^{\prime}_{2}$ if $I_{1}\mathcal{O}(I_{2})=a\mathcal{O}(I_{2})$, and is $I^{\prime}_{2}(\mathcal{O}_{v}:I_{1})^{\diamondsuit}$ if $I_{1}\mathcal{O}(I_{2})$ is a nonprincipal $\mathcal{O}(I_{2})$-ideal.
@@ 3.6. Application to the computation of annihilators in special cases
Since vJ=vb+(n1)vIvJ=vb+(n-1)vI, part 3) of Lemma 2.13 together with Corollary 2.14 shows that ${\mathcal{G}}(vI)={\mathcal{G}}(vJ)$, which implies that $\mathcal{O}(I)=\mathcal{O}(J)$.
@@
2): The value group of $w$ is not discrete, and hence dense, if and only if $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal. If it is discrete and $\gamma$ is its smallest positive element, then $\mathcal{M}=\{b\in K\mid wb\geq\gamma\}$ and $\mathcal{M}^{n}=\{c\in K\mid wc\geq n\gamma\}\subsetneq\mathcal{M}$ since $n\gamma>\gamma$. If it is dense, then for every $b\in\mathcal{M}$ there is $c\in K$ such that $0<nwc<wb$, whence $b\in\mathcal{M}^{n}$; therefore, $\mathcal{M}^{n}\subseteq\mathcal{M}\subseteq\mathcal{M}^{n}$ and consequently, $\mathcal{M}^{n}=\mathcal{M}$.
@@
[…]
@@
2): Assume first that $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}=\mathcal{M}_{v}\,$. Then $a\mathcal{M}/(a\mathcal{M})^{n}$ is nontrivial, whence $a\mathcal{M}\neq(a\mathcal{M})^{n}$, which by part 3) of Lemma 3.20 means that $a\in\mathcal{M}$ or $\mathcal{M}$ is a principal $\mathcal{O}$-ideal. If the latter holds, then by part 1), $a\mathcal{M}:(a\mathcal{M})^{n}=(a\mathcal{M})^{n-1}$. Since $a\mathcal{M}\subseteq\mathcal{M}_{v}\,$, this is equal to $\mathcal{M}_{v}$ if and only if n1=1n-1=1, $\mathcal{M}=\mathcal{M}_{v}\,$, and $va=0$, where we have used parts 1) and 3) of Lemma 3.20. On the other hand, if the latter does not hold, then $a\in\mathcal{M}$ and by part 1), $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}=(a\mathcal{O})^{n-1}\subseteq a\mathcal{O}$. The latter is contained in $\mathcal{M}$ but is not equal to $\mathcal{M}$ since $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal. Thus $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}=a\mathcal{M}:(a\mathcal{M})^{n}\subsetneq\mathcal{M}\subseteq\mathcal{M}_{v}$ and consequently, $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}\neq\mathcal{M}_{v}\,$.